OpenStudy (anonymous):

Use the one-to-one property to solve the given logarithmic equation ln(x+5) + ln(x-6) - ln x = ln 6

OpenStudy (triciaal):

when you add logs multiply the numbers when you subtract logs divide the numbers

OpenStudy (triciaal):

the equal sign means they have the same value

Nnesha (nnesha):

first apply product and property and quotient property \[\huge\rm ln x + \ln y = \ln (x \times y )\] quotient property \[\huge\rm ln x - \ln y = \ln \frac{ x }{ y }\]

OpenStudy (triciaal):